Two cars start off at the same point on a straight highway facing opposite directions. Each car drives for 6 miles, takes a left turn, and drives for 8 miles. How far apart are the two cars?
Answer by Paul(988)

After the car drive and take a left turn they form a right triagnle.
Using pythagoren theorm we can determine how apart the cars are.
6^2+8^2=c^2
36+64=c^2
100=c^2
c=10

Hence, they are 10 miles apart
Paul.